### v1

@@ Results
Bias signatures in collider reasoning: no / weak explaining away and Markov violations In a collider graph, two qualitative signatures sharply distinguish normative causal inference from common human shortcuts. First, explaining away (EA) occurs when evidence for one cause reduces the belief in the alternative cause in the presence of the effect. We quantify the level of explaining away (EA) directly from agents’ raw normalized probability judgments $\mathrm{EA}\!=\!\Pr(C_{1}{=}1\!\mid\!E\!{=}\!1,C_{2}\!{=}\!0)\!-\!\Pr(C_{1}\!{=}\!1\!\mid\!E\!{=}\!1,C_{2}\!{=}\!1)$, where $EA\!\leq\!0$ represents absense of explaining away and the greater EA is, the more explaining away occurs. Qualitatively, explaining away corresponds to a positive slope in Fig.˜3(c). Second, Markov violation captures the violation of the independence of causes in the absence of evidence about the effect: $\mathrm{MV}\!=\!\big|\Pr(C_{1}{=}1\!\mid\!C_{2}{=}1)\!-\!\Pr(C_{1}{=}1\!\mid\!C_{2}{=}0)\big|,\quad\text{with Markov compliance when }\mathrm{MV}\!\approx\!0.$ Qualitatively Markov compliance corresponds to a flat line in Fig.˜3(b). Humans frequently deviate on both: they often show too little or no explaining away (?, ?, ?) meaning evidence for one cause fails to reduce belief in an alternative cause. Humans also have been shown to violate cause-independence (Markov violation), meaning their judgment of one cause is influenced by the presence or absence of an alternative cause (?, ?, ?, ?, ?). Whether LLMs show similar human biases is an open question we address in this work.
LLMs rarely reproduce human collider biases Across agents, explaining away is common and typically much stronger than in humans indicated by the mean and median EA-levels for LLMs in Fig.˜7 falling above the human baseline (pink, $\mathrm{EA}\!\approx\!0.1$). Under overload, many LLMs show reduced EA relative to baseline, with CoT largely restoring EA toward baseline levels.
Most LLMs are Markov compliant; a minority show associative bias like humans. For the Markov condition, the modal pattern is compliance: most LLMs cluster tightly around $\mathrm{MV}\!=0\!$ in Fig.˜8, indicating that one cause does not spuriously change belief in the other whereas for humans this is true. Under overload, Markov violations occur more frequently, albeit still less than in humans.
Q5: Do LLMs Reason Robustly Under Content Manipulations?
Robustness remains highly model-dependent; Gemini-2.5-pro is near-invariant across all 8 experimental conditions and metrics. CoT typically increases robustness.
@@
Discussion
Our results suggest that many contemporary LLMs apply causal reasoning in collider graphs in a comparatively rule-like manner. Humans, by contrast, allocate more mass to unmentioned background factors (lower Background-adjusted causal strength; Fig.˜6), consistent with more open-world interpretations of the prompt. Yet LLMs’ strong rule-following reasoning style risks failure in real-world settings where uncertainty is intrinsic, underscoring the need to better characterize LLM reasoning strategies to guide their safe and effective application. Despite being trained largely on human-generated text, most LLMs do not reproduce human collider biases: they show strong explaining away (EA) and are largely Markov-compliant, whereas humans exhibit low EA and frequent Markov violations. Robustness is strongly model-dependent and shaped by prompting: abstraction and overload shift causal signatures for many models, while CoT often restores more stable behavior; newer, larger models (e.g., Gemini-2.5-pro) are near-invariant across content manipulations, whereas smaller/older models are markedly more sensitive. Our results further show that both human and machine causal judgments can be predicted by—and compressed into—small interpretable causal Bayes nets. More broadly, this illustrates how Bayesian models can provide computational-level targets that help interpret and diagnose the behavior of multi-billion paramter neural networks, exemplifying that Bayesian and neural network models are complementary in understanding intelligence (?, ?).
(a) Gemini-2.5-Pro: tight cluster indicates robust reasoning.
(b) Gemini-2.5-flash-lite: Wide cluster under direct prompting indicates non-robust reasoning; with CoT increasing robustness.
Figure 9: Causal reasoning robustness across experimental conditions for 2 LLMs and 2 metrics discussed. Tight clustering indicates invariance to content manipulations and prompt variations.

### v2

@@ 3.4 Q4: Are LLMs biased like humans?
Bias signatures in collider reasoning: no / weak explaining away and Markov violations In a collider graph, two qualitative signatures sharply distinguish normative causal inference from common human shortcuts. First, explaining away (EA) occurs when evidence for one cause reduces the belief in the alternative cause in the presence of the effect. We quantify the level of explaining away (EA) directly from agents’ raw normalized probability judgments $\mathrm{EA}\!=\!\Pr(C_{1}{=}1\!\mid\!E\!{=}\!1,C_{2}\!{=}\!0)\!-\!\Pr(C_{1}\!{=}\!1\!\mid\!E\!{=}\!1,C_{2}\!{=}\!1)$, where $EA\!\leq\!0$ represents absense of explaining away and the greater EA is, the more explaining away occurs. Qualitatively, explaining away corresponds to a positive slope in Figure 2(c). Second, Markov violation captures the violation of the independence of causes in the absence of evidence about the effect: $\mathrm{MV}\!=\!\big|\Pr(C_{1}{=}1\!\mid\!C_{2}{=}1)\!-\!\Pr(C_{1}{=}1\!\mid\!C_{2}{=}0)\big|,\quad\text{with Markov compliance when }\mathrm{MV}\!\approx\!0.$ Qualitatively Markov compliance corresponds to a flat line in Figure 2(b). Humans frequently deviate on both: they often show too little or no explaining away Fernbach and Rehder (2013); Rottman and Hastie (2014) meaning evidence for one cause fails to reduce belief in an alternative cause. Humans also have been shown to violate cause-independence (Markov violation), meaning their judgment of one cause is influenced by the presence or absence of an alternative cause Ali et al. (2011); Mayrhofer and Waldmann (2016); Park and Sloman (2013); Rehder and Waldmann (2017). Whether LLMs show similar human biases is an open question we address in this work.
@@
LLMs rarely reproduce human collider biases Across agents, explaining away is common and typically much stronger than in humans indicated by the mean and median EA-levels for LLMs in Figure 7 falling above the human baseline (pink, $\mathrm{EA}\!\approx\!0.1$). Under overload, many LLMs show reduced EA relative to baseline, with CoT largely restoring EA toward baseline levels.
Most LLMs are Markov compliant; a minority show associative bias like humans. For the Markov condition, the modal pattern is compliance: most LLMs cluster tightly around $\mathrm{MV}\!=0\!$ in Figure 8, indicating that one cause does not spuriously change belief in the other whereas for humans this is true. Under overload, Markov violations occur more frequently, albeit still less than in humans.
(a) Gemini-2.5-pro reasons most robustly indicated by tight clusters among all 8 experimental conditions, showing content and prompting invariance.
(b) Gemini-2.5-flash-lite shows clear prompting effects indicated by color-separated clusters (Direct , CoT ) with CoT increasing content invariance drastically.
Figure 9: Causal reasoning robustness across experimental conditions for metrics discussed in this work. Tight clustering indicates greater robustness to content manipulations and prompt variations.
3.5 Q5: Do LLMs Reason Robustly Under Content Manipulations?
Robustness remains highly model-dependent; Gemini-2.5-pro is near-invariant across all 8 experimental conditions. CoT typically increases robustness.
@@ 3.5 Q5: Do LLMs Reason Robustly Under Content Manipulations?
4 Discussion
Our results suggest that many contemporary LLMs apply causal reasoning in collider graphs in a comparatively rule-like manner. Humans, by contrast, allocate more mass to unmentioned background factors (lower Background-adjusted causal strength; Figure 6), consistent with more open-world interpretations of the prompt. Yet LLMs’ strong rule-following reasoning style risks failure in real-world settings where uncertainty is intrinsic, underscoring the need to better characterize LLM reasoning strategies to guide their safe and effective application. Despite being trained largely on human-generated text, most LLMs do not reproduce human collider biases: they show strong explaining away (EA) and are largely Markov-compliant, whereas humans exhibit low EA and frequent Markov violations. Robustness is strongly model-dependent and shaped by prompting: abstraction and overload shift causal signatures for many models, while CoT often restores more stable behavior; newer, larger models (e.g., Gemini-2.5-pro) are near-invariant across content manipulations, whereas smaller/older models are markedly more sensitive. Our results further show that both human and machine causal judgments can be predicted by—and compressed into—small interpretable causal Bayes nets. More broadly, this illustrates how Bayesian models can provide computational-level targets that help interpret and diagnose the behavior of multi-billion paramter neural networks, exemplifying that Bayesian and neural network models are complementary in understanding intelligence (Griffiths et al., 2024).